# Work-From-Home and Privacy: What Do Workers Face and What are They Doing About it?

### v1

@@ 2.2. Recruitment
We recruited gender-balanced survey participants using Prolific (Prolific, 2024), a research-oriented online participant recruitment platform. We required participants to be at least 18 years old, reside in the United States, able to read and write in English, and have at least a 90% approval rate on Prolific. We paid participants \$1.00 for completing the screening survey, and \$3.75 for completing the main survey. The survey was published on November 14, 2023, and closed in the same day after we reached 228 responses.
@@ 2.3. Analysis
We analyzed quantitative data using descriptive statistics. To qualitatively analyze the open-ended responses, two researchers used Template Analysis, a style of thematic analysis that combines both inductive and deductive coding, with emphasis on hierarchical coding without specific prescription regarding the number of levels required and what levels represent (King, 2024; Brooks and King, 2014). The first researcher acted as a main coder and the second as a reviewer. Both researchers are trained and have experience in conducting qualitative analysis. Moreover, a single coder is deemed acceptable in the Template Analysis method and in HCI research (King, 2024; McDonald et al., 2019). Both researchers first familiarized themselves with the data by reading through the responses. Next, the first researcher created an initial codebook (the template), and then coded the whole dataset. The first researcher adjusted the codebook during the coding process. After the first researcher finished coding the whole dataset, the second researcher reviewed the coding applied by the first researcher. Both researchers discussed and resolved disagreements.
To test significance we used the Chi-square test ($\chi^{2}$(mch, 2013), given that no more than 20% of the expected frequencies are less than 5 and none of the cells are less than 1 (Kim, 2017). We used an alpha level of .05 for our statistical tests. All tests were computed using the R statistical tool (The R Foundation, 2024). We computed the Chi-square effect size using Cramer’s V ($\alpha_{c}$).
For analysis purposes and throughout the paper, a scenario is considered uncomfortable if participants felt either “very uncomfortable” or “somewhat uncomfortable” when they experienced it.
2.4. Limitations
Prolific crowdworkers are known to be more technically skilled than the average population. However, Prolific has been shown to provide reasonably generalizable results in security and privacy research (Tang et al., 2022). We used self-reported surveys, which are naturally prone to recall bias and social desirability effects. To mitigate recall bias, we asked participants to answer based on their experiences within the last three years; nevertheless, participants might still have trouble accurately remembering incidents within that time frame. We also acknowledge a limitation in the follow-up questions about the scenarios that caused discomfort where we only asked participants about one scenario they reported as the most memorable scenario that made them feel uncomfortable. However, asking participants follow-up questions about multiple scenarios would likely result in survey fatigue that would affect the quality of the data. We also did not ask participants what conference software they use. Different software may have different default settings and features. Moreover, some participants may not be fully aware of their software’s default settings and features. Finally, our sample is limited to US participants. The prevalence of our scenarios and what might be considered uncomfortable may vary based on social norms, which need to be considered when interpreting our results outside the US context.
@@ 3. Results
In Section 3.1, we describe our participants. In Section 3.2, we describe the prevalence of privacy-invasive scenarios reported by participants. In Section 3.3 and Section 3.4 respectively, we provide a summary of the measures participants took to prevent the presented privacy-invasive scenarios from happening (or happening again), and the reasons for not taking measures if they did not take any. Finally, in Section 3.5, we highlight further characteristics of the scenarios that caused discomfort. We refer to participants as P followed by their ID (P_#). Finally, while our analysis is qualitative, we provide frequencies of main themes to offer a sense of their prevalence in our dataset.
@@ 3.1. Participants
We screened 517 participants. Out of those, 244 met the inclusion criteria and were invited to take the main survey, and 228 completed the main survey. We excluded 14 invalid responses. We ended up with 214 completed and valid responses included in our analysis.
@@
Out of the 214 participants in the final dataset, 116 (54.2%) were males, 94 (43.9%) were females, 3 (1.4%) were non-binary, and one preferred not to answer. Their ages ranged from 18 to 74 years, with a majority (69.2%) falling between 25 to 44 years old. In compliance with our screening requirements, all participants were professional workers (full-time or part-time employees, self-employed, freelancers, or business owners), who had been working from home one or more days per week for at least 3 months (at the time of the survey), were not performing their reported jobs through crowdworking platforms such as Prolific and MTurk, [^2] As these differ from the traditional work model and relationships (e.g. with managers and colleagues) that we assumed when we developed our survey’s scenarios. were communicating with others via conferencing tools such as Zoom at least once every few WFH weeks, used a device that has a screen (smartphone, tablet, laptop or PC), were using video conference software (e.g. Zoom) to facilitate WFH. See Table 5 in Appendix A for further demographic details.
@@
The majority of participants reported they had been working from home for a fairly long time, with 79.4% reporting that they had been working from home at least one day per week for at least two years. For all or most of their meetings, 149 (69.6%) share microphone, 102 (47.7%) camera, and 28 (3.7%) screen. Most participants were living with other people (e.g. spouse and children), while only 31 (14.5%) reported that they were living with “no one.” 170 (79.4%) participants said they have a dedicated “space” to work from home. However, when we asked them to describe the space in an open-ended question, only 59 (27.6%) mentioned having a dedicated “office room,” and the rest were working from places other than an office room, such as bedroom (51), secondary bedroom (34), living room/living room-table (35), dining room/dining-table (14), kitchen/kitchen-table (10). See Table 6 in Appendix A for further details about participants’ WFH setting.
@@ 3.2. Prevalence of Privacy-Invasive Scenarios
Our results show that experiencing privacy-invasive scenarios while working from home is common, where 201 (93.9%) participants experienced at least one of the 14 presented scenarios, and 140 (65.4%) felt uncomfortable during their experience for at least one of the scenarios. The average number of scenarios the participants experienced is 5.2, and the average number of scenarios that made the participants feel uncomfortable is 2.2 scenarios. See Figure 5Figure 6, and Figure 7, and. Table 7 in Appendix A for further details and a breakdown of results for the scenarios our participants experienced and how they felt about them.
We first look at the scenarios experienced by the largest number of participants (regardless of whether they caused discomfort or not), aggregated by category. As Figure 2 illustrates, audio scenarios are the most commonly experienced scenarios, with 187/214 (87.4%) participants experiencing at least one audio scenario. The most experienced types of audio scenarios (from most to least experienced) were related to leaked sound or voice of an object, a pet, the subject, an adult, and a child. On the other hand, autonomy scenarios are the least experienced scenarios, with 72/214 (33.6%) participants experiencing at least one autonomy scenario. A breakdown of results for each scenario is presented in Figure 8 in Appendix A.
@@
Second, we focus on the scenarios that caused the most discomfort. As shown in Figure 3, audio scenarios were most often reported as causing discomfort, with discomfort reported by 107/214 (50%) participants for one or more audio scenarios. A breakdown of results for each scenario is presented in  Figure 9 in Appendix A.
@@
Finally, we present results showing which aggregated scenario categories caused the most discomfort to those who experienced them. That is, the percentage of those who experienced discomfort to the total number of people who experienced one or more scenarios from that category. As Figure 4 illustrates, the autonomy scenario category has the highest percentage, with 54/72 (75.0%) of those who experienced at least one autonomy scenario reporting that they felt uncomfortable, suggesting that autonomy scenarios are the most privacy-invasive scenarios when they occur. Audio scenarios come next with 107/187 (57.2%), while video scenarios remain the least reported scenario category that caused discomfort with 55/138 (39.9%) participants. A breakdown of results for each scenario is presented in Figure 10 in Appendix A.
Table 8 in Appendix A provides a breakdown of the numbers summarized in Figure 2Figure 3, and Figure 4.
@@ 3.3. Measures Adoption
At the end of each scenario matrix question (scenario category), we asked participants whether they took any protective measures to prevent the presented scenarios from happening (if they never experienced any) or happening again (if they experienced any). We asked the participants who reported that they took measures to describe the protective measures they took “including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.” In what follows, we summarize the privacy-protective measures adoption rate and types of measures our participants took.
3.3.1. Measures Adoption Rate
As shown in Table 2, our results show that participants who experienced scenarios in a category were more likely to take protective measures. To test the significance of the measures’ adoption rate between those who experienced one or more scenarios of a particular category or never experienced any, we computed the Chi-square test in each scenario category. The difference in measures adoption rate between those who experienced at least one scenario of a category and those who never did is statistically significant in all categories except the audio category (audio: $\chi^{2}$ = 0.989; p-value = 0.3198; $\alpha_{v}$ = 0.0537), (video: $\chi^{2}$ = 5.1425; p-value = 0.023; $\alpha_{v}$ = 0.145), (data: $\chi^{2}$ = 7.142; p-value = 0.007; $\alpha_{v}$ = 0.173), (autonomy: $\chi^{2}$ = 19.866; p-value ¡ 0.001; $\alpha_{v}$ = 0.292).
In addition, whether participants experienced any of the scenarios in a particular category or not, in all scenarios except autonomy scenarios, nearly half of the participants said they adopted measures to prevent audio (57.0%), video (48.6%), and data (44.4%) scenarios; while only 22.4% participants reported that they adopted measures to prevent autonomy scenarios.
@@ 3.3.1. Measures Adoption Rate
Audio 122/214 57.0% 92/214 43.0% 109/187 (58.3%) 78/187 (41.7%) 13/27 (48.1%) 14/27 (51.9%)
Video 104/214 48.6% 110/214 51.4% 75/138 (54.3%) 63/138 (45.7%) 29/76 (38.2%) 47/76 (61.8%)
Data 95/214 44.4% 119/214 55.6% 50/91 (54.9%) 41/91 (45.1%) 45/123 (36.6%) 78/123 (63.4%)
Autonomy 48/214 22.4% 166/214 77.6% 29/72 (40.3%) 43/72 (59.7%) 19/142 (13.4%) 123/142 (86.6%)
@@ 3.3.2. Types of Measures Taken
To prevent audio scenarios, most participants control the microphone. The most mentioned microphone control measures were keeping the microphone muted unless they needed to talk (43) and using mute whenever they needed to prevent unwanted audio from being transmitted (12) as P_220 described: “I try to keep myself muted unless I’m talking. I’m usually pretty good about muting/unmuting.” Few participants (16) explicitly mentioned using software or hardware settings or features to control the microphone. For example, using noise cancellation features built into the conference software (6) or their headset (6) “to only pick up my voice and minimize any outside noise” (P_043), using “push-to-talk” to provide temporal unmute by pressing a key or button (3), or configuring the mute setting to be on by default (2).
@@
In the same vein, to control video scenarios, participants control the camera. The most mentioned camera control was covering the camera (23), e.g. with a tape or slider as P_068 who puts “a Post-It note over my camera lens so that I would have to physically take it off before appearing on camera,” keeping the camera off unless needed (8), and turning the camera off whenever needed (6). Few participants (25) explicitly mentioned using software settings to control their camera, such as the virtual or blurred background features available on most conference software (22), or configuring the camera off as their default setting (3).
@@
For data scenarios, unlike audio and video scenarios, controlling the shared screen is less practical. However, 3 participants mentioned that they limit screen sharing unless necessary: “I try not to share my screen and do what I can to avoid it” (P_097). Few participants (8) mentioned software settings, features, or tools to control what is shown on the screen. For example, 5 mentioned window sharing instead of full-screen sharing as P_076 explained: “ensure that I don’t share my entire screen, but only share the specific program or tab that needs to be shared,” 2 mentioned using private browsing, one mentioned using software that allows to “share a specific section of my screen” (P_074), one mentioned auto-delete the history, and one mentioned using an ad blocker.
@@
Autonomy scenarios are scenarios in which workers are required to continuously share their microphone or camera. Nonetheless, some participants mentioned that they control their microphone or camera as preventative measures for autonomy scenarios, despite the fact that this is prohibited by their employer. For example, participants mentioned turning their microphone off (7) or keeping mute unless needed (4). Similarly, participants mentioned turning their camera off (6), covering their camera (2), or keeping the camera off unless needed (1). For example, P_027 explained: “if i have a true emergency and need to speak to someone - i will simply turn my camera and microphone off for a minute.” This indicates a tension and conflicting situation between the employer’s stated requirements in the scenarios (disallowing workers from turning off their microphone or camera) and what some participants do in such situations as P_195 expressed: “I decided that I am an adult and if I need to leave my desk or mute my microphone it won’t be the end of the world, if my job wants me to be productive as an adult then I need to be treated as an adult.”
@@
Changes to the surrounding space, either physical or virtual space (the latter for data scenarios), were mentioned across all types of scenarios. Changes to the physical space included closing the door, mentioned by 27 to prevent audio scenarios, and by 15 to prevent video scenarios, while 2 participants mentioned it in relation to autonomy scenarios to mitigate the reasons they might want to stop the camera or microphone. Using a door sign to signal they were in a meeting was also mentioned by a few participants in the audio (4), video (1), and autonomy scenarios (2). For example, P_068 explained they “put a ’Please don’t knock’ sign on my door.” To prevent data scenarios, participants mentioned changes to the virtual space that are equivalent to closing the door in the physical space, such as closing or minimizing the browser, apps, or tabs (32). P_050 explained that they “always make sure to close everything before meetings.”
@@
Participants also mentioned the use of a designated or quiet physical space to prevent audio (6), video (9), and autonomy scenarios (1). Changing the place altogether was mentioned by 4 participants in the audio scenarios, such as P_213 who would “leave the house when it’s getting clean (and work elsewhere),” and by 4 in autonomy scenarios, but none of the participants mentioned changing the place in video scenarios. Equivalently, using a designated space or changing the place to prevent data scenarios were expressed through virtual means. For example, 24 participants mentioned using a separate device for work (including two participants who mentioned virtual machines), such as P_091 who explained: “I do not mix my business devices with my personal devices.” Participants also mentioned using a second monitor (6) to “move certain tabs to a different monitor that I don’t share” (P_196), a separate browser (5), and a separate account for work (4).
@@
In limited cases, participants mentioned involving people at work-related meetings or calls. This was mentioned by 3 in relation to audio scenarios, for example: “if there is something going on outside my home, such as construction. I usually just apologize and let them know what is going on and when they can expect the noise to stop” (P_091). 3 participants in autonomy scenarios would inform people at work when they need to leave.
@@
Vigilance was expressed through being mindful of the physical or virtual surroundings, mostly in relation to video, data, and autonomy scenarios. In video scenarios 23 participants mentioned being mindful of the camera position and what it shows as P_077 described: “I make sure that my private living area cannot be seen.” Only 2 participant in audio scenarios mentioned being mindful of their surroundings. In autonomy scenarios, mindfulness was expressed through planning ahead of time and time management, as mentioned by 15. For example, participants repeatedly mentioned comments such as: “I make sure door is closed during meetings and have everything I need” (P_049), “I try to get everything done” (P_043), and “I try to go to the bathroom” (P_024), before the meeting time. Equivalently, in data scenarios, participants mentioned being mindful of what is shown on the screen (20), mindful of their activity on the work device (19), or their activity during the work or meeting time (4).
Vigilance was also expressed through checking the device status. In audio scenarios, checking the microphone status was mentioned by 11 participants, with 6 of them mentioning repeated checks at various frequencies: “double check” (P_112), “multiple times” (P_033), “triple-check” (P_166), and “constantly keep ensuring” (P_185). In video scenarios, checking the camera status was also mentioned but at a lower rate (possibly because if the camera is on, this is immediately reflected on the screen, unlike the mic), where only 3 mentioned it, with only one of those participants mentioned repeated camera checks.
3.4. Not Adopting Measures
At the end of each scenario matrix question (scenario category), we asked participants who reported they did not take any measures to prevent the presented scenarios from happening (if they never experienced any) or happening again (if they experienced any) about the reasons for not taking any measures. As shown earlier in Table 2, our results show that many participants, whether they experienced any of the scenarios in a particular category or never did, said they did not take any measures to prevent the scenarios from happening (or happening again): audio (43.0%), video (51.4%), data (55.6%), and autonomy (77.6%) scenarios.
Participants who said that they did not take any protective measures provided a variety of reasons for not doing so. We identified six prominent themes that encompass most of these reasons: never experienced, oblivious to measures, understanding work environment, nothing can be done, unbothered, and have nothing to hide. In what follows, we elaborate on each theme. Since these themes are universal across all scenario categories, for readability, we refer the reader to Table 3 for themes’ frequencies. Note that the reasons for not taking measures are non-exclusive.
@@ 3.4. Not Adopting Measures
Never Experienced

### v3

@@ 3.2. Recruitment
Our sample size choice (214 participants) was motivated by balancing cost-effectiveness and acceptable statistical power for scientific research (Serdar et al., 2021). A post hoc power analysis using the G*Power statistical tool (Buchner et al., 2024) for the Chi-square ($\chi^{2}$) test family with Effect size 0.3 (medium); $\alpha$ err prob = 0.05; total sample size: 214; Df = 1, gives us a Power (1 - $\beta$ err prob) = 0.99, suggesting that 214 provides sufficient power for our Chi-square statistical tests.
We recruited gender-balanced survey participants using Prolific (Prolific, 2024), a research-oriented online participant recruitment platform. We required participants to be at least 18 years old, reside in the United States, able to read and write in English, and have at least a 90% approval rate in Prolific. We paid participants \$1.00 for completing the screening survey and \$3.75 for completing the main survey. The survey was published on November 14, 2023, and closed in the same day after we reached 228 responses.
@@ 3.3. Analysis
We used mixed methods to analyze our data. We analyzed quantitative data using descriptive statistics. To test significance, we used the Chi-square test ($\chi^{2}$) (McHugh, 2013), given that no more than 20% of the expected frequencies are less than 5 and none of the cells are less than 1 (Kim, 2017). We used an alpha level of .05 for our statistical tests. All tests were computed using the R statistical tool (The R Foundation, 2024). We computed the Chi-square effect size using Cramer’s V ($\alpha_{c}$).
To qualitatively analyze open-ended responses, two researchers used Template Analysis, a style of thematic analysis that combines both inductive and deductive coding, with an emphasis on hierarchical coding without specific prescription regarding the number of levels required and what the levels represent (King, 2024; Brooks and King, 2014). The first researcher acted as the main coder and the second as a reviewer. Both researchers are trained and have experience conducting qualitative analysis. Moreover, a single coder is deemed acceptable in the Template Analysis method and HCI research (King, 2024; McDonald et al., 2019). Both researchers first familiarized themselves with the data by reading the responses. Next, the first researcher created an initial codebook (the template), and then coded the whole dataset. The first researcher updated the codebook during the coding process. After the first researcher finished coding the entire data set, the second researcher reviewed the coding applied by the first researcher. Both researchers discussed and resolved disagreements and adjusted the codebook and the coding throughout the process as needed.
For analysis purposes and throughout the paper, a scenario is considered uncomfortable if the participant reported that they felt either “very uncomfortable” or “somewhat uncomfortable” when they experienced it. We grouped these categories for analysis for reasons deemed appropriate to collapse categories (de Vaus, 2014; Dusen and Nissen, 2020): to make them more relevant to our research problem and to highlight patterns, especially as we combined relevant categories with low frequencies. The “very uncomfortable” category has low frequencies (see Table 9 in Appendix B for details). The scope of our exploratory study does not aim to investigate the severity of discomfort, thus treating the two categories (i.e. “very uncomfortable” and “somewhat uncomfortable”) separately will not add new information to our findings.
@@ 4. Results
We first describe our participants. We then describe the prevalence of privacy-invasive scenarios reported by participants, the measures participants took to prevent the presented privacy-invasive scenarios from happening (or happening again), and the reasons for not taking measures if they did not take any. Finally, we highlight further characteristics of the scenarios that caused discomfort. We refer to the participants as P followed by their ID (P_#). Finally, we provide frequencies of the main qualitative themes to offer a sense of their prevalence in our dataset.
@@ 4.1. Participants
We screened 517 participants. Out of those, 244 met the inclusion criteria and were invited to take the main survey, and 228 completed the main survey. We excluded 14 invalid responses: 5 showed a pattern of very low-quality answers in the open-ended questions indicating inattention; 2 reported they do not share camera, microphone, or screen in any of their meetings in response to screening Q.15–Q.17; 3 duplicate responses; 2 finished the main survey in less than 3 minutes (well below the average time to complete the survey); one response had a technical error; and one reported in the comment at the end of the survey that they misunderstood some questions (our survey purposefully did not have a back button to edit the answers). We ended up with 214 completed and valid responses included in our analysis.
@@
Of the 214 participants in the final dataset, 116 (54.2%) were males, 94 (43.9%) were females, 3 (1.4%) were non-binary, and one preferred not to answer. Their ages ranged from 18 to 74 years, and the majority (69.2%) fell between 25 and 44 years old. In compliance with our screening requirements, all participants were professional workers (full-time or part-time employees, or self-employed, freelancers, or business owners); had been working from home one or more days per week for at least 3 months (at the time of the survey); were not performing their reported jobs through crowdworking platforms such as Prolific and MTurk [^2] These differ from the traditional work model and relationships (e.g. with managers and colleagues) that we assumed when we developed our survey’s scenarios.; were communicating with others via conferencing tools such as Zoom at least once every few WFH weeks; were using a device that has a screen (smartphone, tablet, laptop or PC) and were using video conference software (e.g. Zoom) to facilitate WFH.
Our participants were employed in a wide range of sectors. However, there were a few dominating sectors: “Information and communication Technology” (22.4%), “Financial” (13.6%), “Sales and retail” (12.1%), and “Health” (10.7%). More than a third of the participants (31.3%) reported “Other.” Our participants’ job titles, provided in an open-text entry, covered a wide range of jobs and levels of seniority. For example, the job titles associated with the most reported sector “Information and communication Technology” include: “Administrative Assistant,” “customer service,” “IT Tech,” “Analyst,” “Senior Data Analyst,” “Developer,” “Project Manager,” and “Consultant.” See Table 5, Table 6,Table 7, and Table 8 in Appendix B for more demographic and WFH context details.
@@
Most of our participants reported having been working from home for a fairly long time, with 79.4% reported having been working from home at least one day per week for at least two years. Regarding the frequencies of sharing (turning on) devices (camera, microphone, and screen), our participants reported that, for all or most of their meetings, 149 (69.6%) shared microphone, 102 (47.7%) shared camera, and 28 (13.1%) shared screen. Most participants were living with other people (e.g. spouse, children, parents, roommates, etc.), while only 31 (14.5%) reported that they were living with “no one.” 170 (79.4%) participants said they have a dedicated “space” to work from home; however, when we asked them to describe the space in an open-ended question, only 59 (27.6%) mentioned having a dedicated “office room,” and the rest were working from places other than an office room, such as bedroom (51), secondary bedroom (34), living room/living room-table (35), dining room/dining-table (14), kitchen/kitchen-table (10). See Table 8 in Appendix B for further details about participants’ WFH setting.
@@ 4.2. Prevalence of Privacy-Invasive Scenarios
Our results show that experiencing privacy-invasive scenarios while working from home is common, where 201 (93.9%) participants experienced at least one of the 14 presented scenarios, and 140 (65.4%) felt uncomfortable during their experience for at least one of the scenarios. The average number of scenarios the participants experienced is 5.2, and the average number of scenarios that made the participants feel uncomfortable is 2.2 scenarios. See Table 9 and Figures Figure 5, Figure 6, and Figure 7 in Appendix B for further details and a breakdown of the results for the scenarios our participants experienced and how they felt about them.
We first look at the scenarios experienced by the largest number of participants (regardless of whether they caused discomfort or not), aggregated by category. As Figure 2 illustrates, audio scenarios are the most commonly experienced scenarios, with 187/214 (87.4%) participants experiencing at least one audio scenario. The most experienced types of audio scenarios (from the most experienced to the least experienced) were related to leaked sound or voice of an object, a pet, the subject, an adult, and a child. On the other hand, autonomy scenarios are the least experienced scenarios, with 72/214 (33.6%) participants having experienced at least one autonomy scenario. A breakdown of results for each scenario is presented in Figure 8 in Appendix B.
@@
Second, we focus on the scenarios that caused the most discomfort. As shown in Figure 3, audio scenarios were most often reported as causing discomfort, with discomfort reported by 107/214 (50%) participants for one or more audio scenarios. A breakdown of results for each scenario is presented in Figure 9 in Appendix B.
@@
Finally, we present results showing which aggregated scenario categories caused the most discomfort to those who experienced them. That is, the percentage of those who experienced discomfort to the total number of people who experienced one or more scenarios from that category. As Figure 4 illustrates, the autonomy scenario category has the highest percentage, with 54/72 (75.0%) of those who experienced at least one autonomy scenario reporting that they felt uncomfortable, suggesting that autonomy scenarios are the most privacy-invasive scenarios when they occur. Audio scenarios come next with 107/187 (57.2%), while video scenarios remain the least reported scenario category that caused discomfort with 55/138 (39.9%) participants. A breakdown of the results for each scenario is presented in Figure 10 in Appendix B.
Table 10 in Appendix B provides a breakdown of the numbers summarized in Figure 2Figure 3, and Figure 4.
@@ 4.3. Measures Adoption
At the end of each scenario matrix question (scenario category), we asked the participants if they took any protective measures to prevent the presented scenarios from happening (if they never experienced any) or happening again (if they experienced any). We asked the participants who reported that they took measures to describe the protective measures they took “including any new software you started using, changes to software settings, changes to your physical workspace, or anything else.” In what follows, we summarize the adoption rate of privacy protection measures and the types of measures our participants took.
4.3.1. Measures’ Adoption Rate
As shown in Table 2, our results show that participants who experienced scenarios in a category were more likely to take protective measures. To test the significance of the measures’ adoption rate between those who experienced one or more scenarios of a particular category or never experienced any, we computed the Chi-square test in each scenario category. The difference in measures’ adoption rate between those who experienced at least one scenario of a category and those who never did is statistically significant in all categories except the audio category (audio: $\chi^{2}$ = 0.989; p-value = 0.3198; $\alpha_{v}$ = 0.0537), (video: $\chi^{2}$ = 5.1425; p-value = 0.023; $\alpha_{v}$ = 0.145), (data: $\chi^{2}$ = 7.142; p-value = 0.007; $\alpha_{v}$ = 0.173), (autonomy: $\chi^{2}$ = 19.866; p-value $<$ 0.001; $\alpha_{v}$ = 0.292).
Furthermore, whether the participants experienced any of the scenarios in a particular category or not, in all scenarios except autonomy scenarios, nearly half of the participants said that they adopted measures to prevent audio (57.0%), video (48.6%), and data (44.4%) scenarios; while only 22.4% of participants reported that they adopted measures to prevent autonomy scenarios.
@@ 4.3.1. Measures’ Adoption Rate
Audio 122/214 (57.0%) 92/214 (43.0%) 109/187 (58.3%) 78/187 (41.7%) 13/27 (48.1%) 14/27 (51.9%)
Video 104/214 (48.6%) 110/214 (51.4%) 75/138 (54.3%) 63/138 (45.7%) 29/76 (38.2%) 47/76 (61.8%)
Data 95/214 (44.4%) 119/214 (55.6%) 50/91 (54.9%) 41/91 (45.1%) 45/123 (36.6%) 78/123 (63.4%)
Autonomy 48/214 (22.4%) 166/214 (77.6%) 29/72 (40.3%) 43/72 (59.7%) 19/142 (13.4%) 123/142 (86.6%)
@@ 4.3.2. Types of Measures Taken
To prevent audio scenarios, most of the participants control the microphone. The most mentioned microphone control measures were keeping the microphone muted unless they needed to talk (43) and using mute whenever they needed to prevent unwanted audio from being transmitted (12) as described by P_220: “I try to keep myself muted unless I’m talking. I’m usually pretty good about muting/unmuting.” Few participants (16) explicitly mentioned using software or hardware settings or features to control the microphone. For example, using noise cancelation features built into conference software (6) or headset (6) “to only pick up my voice and minimize any outside noise” (P_043), using “push-to-talk” to provide temporal unmute by pressing a key or button (3), or configuring the mute setting to be on by default (2).
@@
In the same vein, to control video scenarios, participants control the camera. The most mentioned camera control was covering the camera (23), e.g. with a tape or slider as P_068 who places “a Post-It note over my camera lens so that I would have to physically take it off before appearing on camera,” keeping the camera off unless needed (8) and turning off the camera whenever needed (6). Few participants (25) explicitly mentioned using software settings to control their camera, such as the virtual or blurred background features available in most conference software (22), or configuring the camera off as their default setting (3).
@@
For data scenarios, unlike audio and video scenarios, controlling the shared screen is less practical. However, 3 participants mentioned that they limit screen sharing unless necessary: “I try not to share my screen and do what I can to avoid it” (P_097). Few participants (8) mentioned software settings, features, or tools to control what is shown on the screen. For example, 5 mentioned window sharing instead of full-screen sharing as P_076 explained: “ensure that I don’t share my entire screen, but only share the specific program or tab that needs to be shared,” 2 mentioned using private browsing, one mentioned using software that allows to “share a specific section of my screen” (P_074), one mentioned auto-deleting the history, and one mentioned using an ad blocker.
@@
Autonomy scenarios are scenarios in which workers are required to continuously share their microphone or camera. Nonetheless, some participants mentioned that they control their microphone or camera as preventative measures for autonomy scenarios, even though this is prohibited by their employer. For example, participants mentioned turning their microphone off (7) or keeping mute unless needed (4). Similarly, the participants mentioned turning off their camera (6), covering their camera (2), or keeping the camera off unless needed (1). For example, P_027 explained: “if i have a true emergency and need to speak to someone - i will simply turn my camera and microphone off for a minute.” This indicates a tension and conflicting situation between the employer’s stated requirements in the scenarios (disallowing workers from turning off their microphone or camera) and what some participants do in such situations as P_195 expressed: “I decided that I am an adult and if I need to leave my desk or mute my microphone it won’t be the end of the world, if my job wants me to be productive as an adult then I need to be treated as an adult.”
@@
Changes to the surrounding space, either physical or virtual space (the latter for data scenarios), were mentioned across all types of scenarios. Changes in physical space included closing the door, mentioned by 27 to prevent audio scenarios, and by 15 to prevent video scenarios, while 2 participants mentioned it in relation to autonomy scenarios to mitigate the reasons they might want to stop the camera or microphone. Using a door sign to signal that they were in a meeting was also mentioned by a few participants in the audio (4), video (1), and autonomy scenarios (2). For example, P_068 explained that they “put a ‘Please don’t knock’ sign on my door.” To prevent data scenarios, participants mentioned changes to the virtual space that are equivalent to closing the door in the physical space, such as closing or minimizing the browser, apps, or tabs (32). P_050 explained that they “always make sure to close everything before meetings.”
@@
The participants also mentioned the use of a designated or quiet physical space to prevent audio (6), video (9), and autonomy scenarios (1). Changing the place altogether was mentioned by 4 participants in the audio scenarios such as P_213 who would “leave the house when it’s getting clean (and work elsewhere),” and by 4 in autonomy scenarios, but none of the participants mentioned changing the place in video scenarios. Equivalently, using a designated space or changing the place to prevent data scenarios were expressed through virtual means. For example, 24 participants mentioned using a separate device for work (including two participants who mentioned virtual machines) such as P_091 who explained: “I do not mix my business devices with my personal devices.” Participants also mentioned using a second monitor (6) to “move certain tabs to a different monitor that I don’t share” (P_196), a separate browser (5), and a separate account for work (4).
@@
In limited cases, participants mentioned involving people at work-related meetings or calls. This was mentioned by 3 in relation to audio scenarios, for example: “if there is something going on outside my home such as construction. I usually just apologize and let them know what is going on and when they can expect the noise to stop” (P_091). 3 participants in autonomy scenarios would inform people at work when they need to leave.
@@
Vigilance was expressed through being mindful of the physical or virtual surroundings, mostly in relation to video, data, and autonomy scenarios. In video scenarios 23 participants mentioned being mindful of the camera position and what it shows as P_077 described: “I make sure that my private living area cannot be seen.” Only 2 participants in audio scenarios mentioned being mindful of their surroundings. In autonomy scenarios, mindfulness was expressed through planning ahead of time and time management, as mentioned by 15 participants. For example, the participants repeatedly mentioned comments such as “I make sure door is closed during meetings and have everything I need” (P_049), “I try to get everything done” (P_043), and “I try to go to the bathroom” (P_024), before the meeting time. Equivalently, in data scenarios, participants mentioned being mindful of what is shown on the screen (20), mindful of their activity on the work device (19), or their activity during the work or meeting time (4).
Vigilance was also expressed by checking the device status. In audio scenarios, checking the microphone status was mentioned by 11 participants, with 6 of them mentioning repeated checks at various frequencies: “double check” (P_112), “multiple times” (P_033), “triple-check” (P_166), and “constantly keep ensuring” (P_185). In video scenarios, checking the camera status was also mentioned, but at a lower rate (possibly because if the camera is on, this is immediately reflected on the screen, unlike the mic), where only 3 mentioned it, with only one of those participants mentioning repeated camera checks.
@@
4.4. Not Adopting Measures
At the end of each scenario matrix question (scenario category), we asked participants who reported that they did not take any measures to prevent the presented scenarios from happening (if they never experienced any) or happening again (if they experienced any) about the reasons for not taking any measures.
As shown earlier in Table 2, our results show that many participants, whether they experienced any of the scenarios in a particular category or never did, said they did not take any measures to prevent the scenarios from happening (or happening again): audio (43.0%), video (51.4%), data (55.6%), and autonomy (77.6%) scenarios.
Participants who said they did not take any protective measures provided a variety of reasons for not doing so. We identified six prominent themes that encompass most of these reasons: never experienced, oblivious to measures, understanding work environment, nothing can be done, unbothered, and have nothing to hide. In what follows, we elaborate on each theme.
Since these themes are universal across all scenario categories, for readability, we refer the reader to Table 3 for themes’ frequencies. Note that the reasons for not taking measures are non-exclusive.
Never experienced.